answersLogoWhite

0

The answer will depend on

  • the deck of cards that you draw from: a pinochle deck, a deck with jokers, etc;
  • how many cards are drawn,
  • if more than one card is drawn, whether or not a drawn card is replaced before the next one is drawn.
User Avatar

Wiki User

∙ 12y ago

What else can I help you with?

Continue Learning about Statistics

What is the probability of drawing a king and a diamond from a deck of playing cards?

The probability of drawing a king is 4:52The probability of drawing a diamond is 13:52 (or 1:4)The probability of drawing a king (0.07692...) then replacing that king into the deck then drawing a diamond is 0.019230769.If you leave the king out, the probability will be slightly greater (4/52) * (13/51)Unless the king you left out of the deck was a king of diamonds, in which case, the probability would be (4/52) * (12/51)


What is the probability of drawing a red king from a deck of 52?

Probability of getting a King is 4/52 = 1/13. Having drawn the King, for it to be a Red King, the probability is 2/4 = 1/2. Hence the conditional probability is (1/13)x(1/2) = 1/26


What is the probability of choosing a queen or a king from a standard deck of cards?

The probability of drawing a queen or king, in a single randomly drawn card, is 2/13. The probability of drawing one when you draw 45 cards without replacement is 1. The probability of choosing has nothing t do with the probability of drawing the card. I can choose a king but fail to find one!


What is the probability of drawing 2 kings from a standard deck of cards with replacement?

This problem is the type of the probability of A and the probability of B. These events are independent. P(A) and P(B) = P(A) * P(B). In this case these two probabilities are equal; the probability of a king is 4/52. So, the probability of draw king, replace and draw king is 4/52 * 4/52 = 0.00592.


What is probability of drawing a king or queen from a deck of cards?

There are 4 kings and 4 queens in a deck of 52 cards. The chance of drawing a king is 4 in 52 (or .077 in probability terms). The probability to draw either a king or a queen will be twice as high (.154)

Related Questions

Did King Henry invented fractions?

Who created fractions first**,Egyptians** or King Henry


What is the probability of drawing a king and a diamond from a deck of playing cards?

The probability of drawing a king is 4:52The probability of drawing a diamond is 13:52 (or 1:4)The probability of drawing a king (0.07692...) then replacing that king into the deck then drawing a diamond is 0.019230769.If you leave the king out, the probability will be slightly greater (4/52) * (13/51)Unless the king you left out of the deck was a king of diamonds, in which case, the probability would be (4/52) * (12/51)


What is the probability of drawing a red king from a deck of 52?

Probability of getting a King is 4/52 = 1/13. Having drawn the King, for it to be a Red King, the probability is 2/4 = 1/2. Hence the conditional probability is (1/13)x(1/2) = 1/26


What is the probability of not getting a king from a deck of 52 cards?

The probability that a standard deck of 52 cards does not contain a king is 0.


What is the probability of getting a king?

What part of the world are you in? If you're talking about the united states replacing the president with a king, that is impossible. There is no probability of the united states getting a king.


What is the probability of choosing a queen or a king from a standard deck of cards?

The probability of drawing a queen or king, in a single randomly drawn card, is 2/13. The probability of drawing one when you draw 45 cards without replacement is 1. The probability of choosing has nothing t do with the probability of drawing the card. I can choose a king but fail to find one!


What is the probability of drawing 2 kings from a standard deck of cards with replacement?

This problem is the type of the probability of A and the probability of B. These events are independent. P(A) and P(B) = P(A) * P(B). In this case these two probabilities are equal; the probability of a king is 4/52. So, the probability of draw king, replace and draw king is 4/52 * 4/52 = 0.00592.


What is the probability of choseing a king or king from a deck of 52?

The probability of getting a king in a one card draw from a deck of 52 cards is .077 or 1 chance in 13. The probability of getting a king and one other specific denomination, such as a queen, is .154, or one chance in between 6 and 7.


What is probability of drawing a king or queen from a deck of cards?

There are 4 kings and 4 queens in a deck of 52 cards. The chance of drawing a king is 4 in 52 (or .077 in probability terms). The probability to draw either a king or a queen will be twice as high (.154)


What is the probability of choosing a king from a standard deck of playing cards?

The probability is 1/13 of drawing a king in one draw from a standard deck with no jokers.


What is the probability of pulling out one card and it being between an 8 and a king?

The probability is 4/13.


The king comes from a family of three children What is the probability that he has a brother?

The probability of a boy (male) is equal to the probability of a girl (female) which equals 1/2. The king is a male. So, we need the probability of a male and a male which is 1/2 * 1/2 = 1/4.

Trending Questions
Does intervals go on the x axis of a bar graph? Sample letter stating you know someone for a long time? How can stock price react randomly to events that have more than 50 percent probability of happening? Why does a 115 pound person drinking 3 beers get drunker than a 155 pound person drinking the same amount? A concert organizer recorded the number of people in different age groups who attended a concert. She then created the following histogram Histogram with title Concert Attendance horizontal axis label? What is does overstating something in order to get some effect? How many ways can five differently colored marbles be arrange in a row? What is average weight of 14 year old? What happens to interval when level of confidence is increased? Here is the histogram of a data distribution The classes all consist of just one number the class width is 1 Which of the following numbers is closest to the mean of this distribution? How do you calculate budget variance? What is the different of straight line graph and broken line graph? What should a 6 foot 66 year old male weigh? What are the odds of pregnancy if the man has a vasectomy? What is a stochastic error? What does a coefficient of correlation of 0.70 infer? What is Rejected Hypothesis? On a calculator what does the HYP button do? You roll a dice once what is the probability that you roll an odd number? If the probability that it will snow tomorrow is 43 what is the the probability that it will not snow?